OpenStudy (anonymous):

Which group of points are all on the line y = -1/2x - 7?

OpenStudy (radar):

Just substitute for x and y using the points and find out which is true.

OpenStudy (anonymous):

subsitute it with what

OpenStudy (radar):

for example if the Point P is (0,1) then test 1 = (-1/2)(0) -7 1=0-7 1=-7 Not true!

OpenStudy (radar):

Substitute the points in question which for some reason you have not given us.

OpenStudy (radar):

Do you understand? Your question is incomplete you need to show the points if you have a grid showing points you must provide it so your question can be answered.

OpenStudy (radar):

However, I have provided a means for you to answer your question. Good luck with it.

OpenStudy (anonymous):

oh ok thanks